Find an equation of a sphere with radius r and center c(h,k,l)
1
Expert's answer
2020-10-04T13:16:02-0400

By defenition of the sphere, the equation is

"(x-c_x)^2+(y-c_y)^2+(z-c_z)^2=r^2," or

"(x-h)^2+(y-k)^2+(z-l)^2=r^2."


Answer: "(x-h)^2+(y-k)^2+(z-l)^2=r^2."